Prev Next

SW&IT Meeting Minutes 2021-02-15

Attendees: David, Igor, Matthieu, Paul, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

20210208
LR20210208 - Bec, Matthieu 32.0
LR20210208 - Edmeades, David 40.0
LR20210208 - Sola, Igor 0.0
LR20210208 - Grenz, Paul 16.0
LR20210208 - Hooper, Stephen 80.0
LR20210208 - Zhang, Yang 40.0
LR20210208 - Zhang, Xianyu 32.0

News from management

  • ATT AORR review

Previous Action Items

Owner Action Due Date Completion Date

Open topics

Team reports

MB

  • Last week
    • ATT S/W slides complete, AORR review
    • HBS plc upgrade discussion
    • AGW hands-on with JP
    • Web student review candidates
    • Off Friday
  • This week
    • SHARK NIR RFAs
    • review candidates

YZ

  • Previous:
    • Spent most time on OSCO software. Now completed all components (realtime computation/database design/wsgi query/polymer web etc.).
    • Also completed all OSCO image analysis since July 01, 2020.
    • Some discussion on TMS software revision.
  • Next:
    • Work with Matthieu to install and configure the live OSCO system in DMS.
    • Work on TMS software revision and possibly scheduling another test.
    • If have time, work on some misc items (TCS doc revision, MODS 64bit stuff, etc.).

DE

  • Power stuff, thanks Stephen
  • Last-minute VPN account request--can we emphasize that the observers need to give us notice?
  • Storage doc work
  • Synology work
    • Ordered used unit, arrival ~20 Feb
    • Syno offered to RMA ours with new, trying to work out how to get an advance RMA done within purchasing rules
  • Talking to SHARK-vis team about switch gear; no response to first questions yet.
  • Need to touch backup machines after power shenanigans.
  • Second round vaccination on Monday 22 Feb, blocking out 22 and 23 in case of immune response

XZ

  • Last week
    • ATT AdSec and SW slides
    • AORR meeting
    • Review SOUL issues
    • Friday off
  • This week
    • Work on the SOUL issues
    • Merge to SOUL to Master
    • AO support

SH

  • Steward power stuff
    • info jobs dying
    • problems with ovirt
  • DIMM
    • LeRoy has materials
  • Meetings -- power, strategic stuff.
  • M3 stuff for Juan Pablo (tornado routing)
  • Powered off archive no complaints

next:

  • finish up Steward power.
  • Mountain ovirt reinstall on "vm2". For some reason it is pathetically slow.
  • Finish robs2
  • Update network and DNS configuration in Git.
  • Power off core-b (mainly LeRoy)
  • DIMM ?

PSG

  • Last week
    • Adding code to OSS to support the ARGOS swing arm.
    • Off Friday.
  • This week
    • Talking to Yang about setting up virtual machine on tcs-test.
    • Compiling and testing new version of OSS.

IS

Last week:
  • OT
  • LUCI
    • IT #7493 Aladin slow subtraction. Doing the subtraction outside Aladin
    • IT #8231- Timeout on GEIRS during readout

This week:
  • OT XSLT test cases
  • PIT new release work
  • LUCI
    • IT #7493 Aladin slow subtraction
Topic revision: r9 - 15 Feb 2021, IgorSola
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback